OpenStudy (anonymous):

9. Ho of ethyl alcohol is -1236 kJ. How much energy can be produced by burning 220g of ethyl alcohol in excess oxygen? C2H6O(l) + 3O2(g)  2CO2 (g) + 3H2O (g)

OpenStudy (cuanchi):

that value of delta H (-1236kJ) is the energy produced by one mole of ethyl alcohol. You have to calculate how many moles of ethyl alcohol you have in 220g ( n=m/MM) and multiply the molar delta H by the number of moles -1236kJ x n = ?

OpenStudy (anonymous):

I'm not entirely sure how to do that my teacher wasn't exactly the best so this stuff really confuses me

OpenStudy (cuanchi):

If you put one log in the fireplace you will get some amount of heat, if you put 2 logs you will have more heat (two times more) if you put 5 or 10 logs you will have 5 or 10 times more heat. The same here, if you put one mole of alcohol you will get -1236 kJ if you put any other amount you will have a proportional amount of heat. The mass of alcohol is given in g you have to convert the grams to moles and then (when you know how many moles, or logs in your fireplace) multiply that by the amount of heat that one mole produce.

OpenStudy (anonymous):

oh ok got it so -1236kj x 220g or wait no because n = m/MM so how do I know or find m/MM to find n? ( sorry I'm so bad at this)

OpenStudy (cuanchi):

n= m/MM n= number of moles m= mass MM= molecular mass you have to calculate the MM of the ethyl alcohol based in its formula and the atomic mas of the elements that compose it. then divide the mass (220g by the MM) to get the number of moles then multiply the delta H by the number of moles

OpenStudy (anonymous):

mmk I think I got it, thank you =)